# Heterochromatic Geometric Transversals of Convex sets

### v4

@@ 2 Definitions and notations
  • $\mathbb{S}^{d-1}$ denotes the unit sphere in $\mathbb{R}^{d}$ with center at the origin $O$.
  • For any $x\in\mathbb{R}^{d}\setminus\{O\}$, $\hat{x}$ denotes the central projection of $x$ on $\mathbb{S}^{d-1}$, that is, $\hat{x}:=\frac{x}{||x||}$.
@@
  • An unbounded cone $\mathcal{C}$ in $\mathbb{R}^{d}$ with vertex at the origin $O$ is a convex set such that for any $u,v\in\mathcal{C}$ and $\alpha,\beta\geq 0$, we have $\alpha u+\beta v\in\mathcal{C}$. The axis of $\mathcal{C}$ is a straight line $\ell$ passing through the vertex such that $\mathcal{C}$ has a circular symmetry about $\ell$. If we translate $\mathcal{C}$ at a point $z\in\ell$, we get the unbounded cone $\mathcal{C}^{\prime}=\{z+c\;|\;c\in\mathcal{C}\}$.
@@
3 Heterochromatic $(\aleph_{0},\cdot)$-theorem for fat convex sets
In this section, we will first prove Theorem 1.10, and then Proposition 1.11 and Theorem 1.12.
Let us first prove a Lemma that we will be using quite frequently. Let $\mathcal{B}$ and $\mathcal{S}$ be collections of nonempty sets in $\mathbb{R}^{d}$. For any $\mathcal{F}\subseteq\mathcal{B}$, we say $\mathcal{F}$ is finitely pierceable by $\mathcal{S}$ if there exists a finite set $S\subseteq\mathcal{S}$ such that for all $B\in\mathcal{F}$, we have $\bigcup_{A\in S}A\cap B\neq\emptyset$. A sequence $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ with $\mathcal{F}_{n}\subseteq\mathcal{B}$ $\forall n\in\mathbb{N}$ is said to have the heterochromatic $(\aleph_{0},q)$-property with respect to $\mathcal{S}$ if for every heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$ chosen from $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$, we have a set $A\in\mathcal{S}$ such that there are $q$ distinct $B_{n}$’s that have a nonempty intersection with $A$. We say that the heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ with respect to $\mathcal{S}$ holds if for any $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ with $\mathcal{F}_{n}\subseteq\mathcal{B}$ satisfying the heterochromatic $(\aleph_{0},q)$-property w.r.t. $\mathcal{S}$, there exists an $n\in\mathbb{N}$ for which $\mathcal{F}_{n}$ is finitely pierceable by $\mathcal{S}$. We say that the infinite heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ w.r.t. $\mathcal{S}$ holds if for any $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ with $\mathcal{F}_{n}\subseteq\mathcal{B}$ satisfying the heterochromatic $(\aleph_{0},q)$-property w.r.t. $\mathcal{S}$, all but finitely many $\mathcal{F}_{n}$’s are finitely pierceable by $\mathcal{S}$.
@@ 3 Heterochromatic $(\aleph_{0},\cdot)$-theorem for fat convex sets
The heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ w.r.t. $\mathcal{S}$ implies the infinite heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ w.r.t. $\mathcal{S}$.
@@
Let the heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ w.r.t. $\mathcal{S}$ hold. Let $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$, $\mathcal{F}_{n}\subseteq\mathcal{B}$, be a sequence that satisfies the heterochromatic $(\aleph_{0},q)$-property w.r.t. $\mathcal{S}$. If possible, let $\{\mathcal{F}_{n_{i}}\}_{i\in\mathbb{N}}$ where $\{n_{i}\}_{i\in\mathbb{N}}$ is a strictly increasing sequence in $\mathbb{N}$ be such that no $\mathcal{F}_{n_{i}}$ is finitely pierceable by $\mathcal{S}$. As $\{\mathcal{F}_{n_{i}}\}_{i\in\mathbb{N}}$ also satisfies the heterochromatic $(\aleph_{0},q)$-property w.r.t. $\mathcal{S}$, there is an $i\in\mathbb{N}$ for which $\mathcal{F}_{n_{i}}$ is finitely pierceable by $\mathcal{S}$. But this is a contradiction, and therefore, the result follows. ∎

### v5

@@ 2 Definitions and notations
  • We will denote by $\mathcal{O}$ the origin in $\mathbb{R}^{d}$.
  • $\mathbb{S}^{d-1}$ denotes the unit sphere in $\mathbb{R}^{d}$ with center at the origin $\mathcal{O}$.
  • For any $x\in\mathbb{R}^{d}\setminus\{\mathcal{O}\}$, $\hat{x}$ denotes the central projection of $x$ on $\mathbb{S}^{d-1}$, that is, $\hat{x}:=\frac{x}{||x||}$.
@@
  • An unbounded cone $\mathcal{C}$ in $\mathbb{R}^{d}$ with vertex at the origin $\mathcal{O}$ is a convex set such that for any $u,v\in\mathcal{C}$ and $\alpha,\beta\geq 0$, we have $\alpha u+\beta v\in\mathcal{C}$. The axis of $\mathcal{C}$ is a straight line $\ell$ passing through the vertex such that $\mathcal{C}$ has a circular symmetry about $\ell$. If we translate $\mathcal{C}$ at a point $z\in\ell$, we get the unbounded cone $\mathcal{C}^{\prime}=\{z+c\;|\;c\in\mathcal{C}\}$.
@@
  • $k$-set: We say $\mathcal{F}$ is a $k$-set if $\mathcal{F}$ is a collection of compact convex sets in $\mathbb{R}^{d}$ which does not have a finite $k$-flat transversal. We say $A\subseteq\mathbb{R}^{d}$ contains a $k$-set $\mathcal{B}$ or $\mathcal{B}$ is contained in $A$ if $\forall B\in\mathcal{B}$, $B\subseteq A$.
[…]
  • For any $S\in\mathcal{F}$ where $\mathcal{F}$ is a collection of sets in $\mathbb{R}^{d}$, $B_{S}$ denotes a largest ball inscribed into $S$ and $x_{S}$ denotes the centre of $B_{S}$.
  • For any $S\in\mathcal{F}$ where $\mathcal{F}$ is a collection of compact sets in $\mathbb{R}^{d}$, $r_{S}$ and $R_{S}$ denote the radius of the largest closed ball contained in $S$ and the radius of the smallest closed ball containing $S$ respectively.
  • For a family $\mathcal{F}$ of sets in $\mathbb{R}^{d}$ and any two positive numbers $r,R\in\mathbb{R}^{+}$, $\mathcal{F}_{r,R}:=\{S\in\mathcal{F}\mid r_{S}\geq r\text{ and }R_{S}\leq R\}$, $\mathcal{F}_{<r}:=\{S\in\mathcal{F}\mid r_{S}<r\}$ and $\mathcal{F}_{>R}:=\{S\in\mathcal{F}\mid R_{S}>R\}$.
3 Convex sets with bounded condition number
In this section we prove Theorem 1.13 in the following three steps:
  • first we show that heterochromatic $(\aleph_{0},.)$-theorem implies the existence of a finite hitting set,
  • next we establish that monochromatic $(\aleph_{0},k+2)$-theorem for compact convex bodies and $k$-flats implies Theorem 1.13, and
  • finally we prove the monochromatic version of Theorem 1.13.
3.1 Heterochromatic $(\aleph_{0},.)$-theorem implies the existence of a finite hitting set
In this section, we shall show that the heterochromatic $(\aleph_{0},q)$-theorem with $q\in\mathbb{N}$ implies the existence of a finite hitting set, regardless of the types of sets chosen from $\mathbb{R}^{d}$. Let us first prove a Lemma. Let $\mathcal{B}$ and $\mathcal{S}$ be collections of nonempty sets in $\mathbb{R}^{d}$. For any $\mathcal{F}\subseteq\mathcal{B}$, we say $\mathcal{F}$ is finitely pierceable by $\mathcal{S}$ if there exists a finite set $S\subseteq\mathcal{S}$ such that for all $B\in\mathcal{F}$, we have $\bigcup_{A\in S}A\cap B\neq\emptyset$. A sequence $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ with $\mathcal{F}_{n}\subseteq\mathcal{B}$ $\forall n\in\mathbb{N}$ is said to have the heterochromatic $(\aleph_{0},q)$-property with respect to $\mathcal{S}$ if for every heterochromatic sequence $\{B_{n}\}_{n\in\mathbb{N}}$ chosen from $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$, we have a set $A\in\mathcal{S}$ such that there are $q$ distinct $B_{n}$’s for which $B_{n}\cap A\neq\emptyset$. We say that the heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ with respect to $\mathcal{S}$ holds if for any $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ with $\mathcal{F}_{n}\subseteq\mathcal{B}$ satisfying the heterochromatic $(\aleph_{0},q)$-property with respect to $\mathcal{S}$, there exists an $n\in\mathbb{N}$ for which $\mathcal{F}_{n}$ is finitely pierceable by $\mathcal{S}$. We say that the infinite heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ with respect to $\mathcal{S}$ holds if for any $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$ with $\mathcal{F}_{n}\subseteq\mathcal{B}$ satisfying the heterochromatic $(\aleph_{0},q)$-property with respect to $\mathcal{S}$, all but finitely many $\mathcal{F}_{n}$’s are finitely pierceable by $\mathcal{S}$.
@@ 3 Convex sets with bounded condition number
The heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ with respect to $\mathcal{S}$ implies the infinite heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ with respect to $\mathcal{S}$.
@@
Let the heterochromatic $(\aleph_{0},q)$-theorem for $\mathcal{B}$ with respect to $\mathcal{S}$ hold. Let $\{\mathcal{F}_{n}\}_{n\in\mathbb{N}}$, $\mathcal{F}_{n}\subseteq\mathcal{B}$, be a sequence that satisfies the heterochromatic $(\aleph_{0},q)$-property with respect to $\mathcal{S}$. If possible, let $\{\mathcal{F}_{n_{i}}\}_{i\in\mathbb{N}}$ where $\{n_{i}\}_{i\in\mathbb{N}}$ is a strictly increasing sequence in $\mathbb{N}$ be such that no $\mathcal{F}_{n_{i}}$ is finitely pierceable by $\mathcal{S}$. As $\{\mathcal{F}_{n_{i}}\}_{i\in\mathbb{N}}$ also satisfies the heterochromatic $(\aleph_{0},q)$-property with respect to $\mathcal{S}$, there is an $i\in\mathbb{N}$ for which $\mathcal{F}_{n_{i}}$ is finitely pierceable by $\mathcal{S}$. But this is a contradiction, and therefore, the result follows. ∎